OpenStudy (anonymous):

LIMIT:

OpenStudy (anonymous):

\[\lim_{x \rightarrow-4} [\sqrt{x^2+9} - 5] / x+4\]

OpenStudy (goformit100):

With due respect Welcome to Open Study ............. Sir/Ma'am @viemer

zepdrix (zepdrix):

\[\Large\rm \lim_{x\to -4}\frac{\sqrt{x^2+9}-5}{x+4}\color{royalblue}{\left(\frac{\sqrt{x^2+9}+5}{\sqrt{x^2+9}+5}\right)}\]What have you tried? This is probably a good place to start. Multiply the top and bottom by the `conjugate` of the top.

OpenStudy (anonymous):

\[x-4 / \sqrt{x^2+9}+5\]

OpenStudy (anonymous):

that's what I got... is it right? Does it end there?

zepdrix (zepdrix):

Hmm that looks good so far. Remember that (x+4) in the denominator was causing a problem? Whenever you are able to get rid of a problem like that, try plugging your limit in. See if it works at this point.

OpenStudy (anonymous):

my final answer is -4/5

zepdrix (zepdrix):

Yay good job \c:/

OpenStudy (anonymous):

thanks ^^